a toy rocket is launched upward from a height of 1.3 m above the ground. The height of the rocket, in meters, after t seconds is modeled by the equation h=-4.9t^2+51t+1.3 What is the maximum height and when does it reach that height?

Answer by Alan3354(69443)

h=-4.9t^2+51t+1.3
--------
Max ht is the vertex at t = -b/2a
t = -51/-9.8 = 255/49
-----
hmax = -4.9(255/49)^2 + 51*255/49 + 1.3
=~ 134 meters
----------------------
PS That's not a rocket. Rockets have thrust.
It's a projectile.